The pressure of a gas is raised from 27°C to 927°C. The root mean square speed is



              


(b)    Remain the same 


(c)    Gets halved


(d)    Get doubled


 

Grade:10

2 Answers

Simran Bhatia
10 years ago

(d)

crms proportional to sqrt(T) 

As temperature increases from 300 K to 1200 K that is four times, so, crms will be doubled.

MADHAN
5 years ago
rms velocity =√3rt/m
since r,m are constants 
so we consider,
v1/v2=√t1/√t2
         =√300/√1200
         =1/2
hence they are doubled